OpenStudy (anonymous):

A street light is at the top of a 19 ft tall pole. A woman 6 ft tall walks away from the pole with a speed of 5 ft/sec along a straight path. How fast is the tip of her shadow moving when she is 40 ft from the base of the pole? Rate =

myininaya (myininaya):

did you try drawing a traingle. You have a triangle inside a triangle. So you should set up the equation (x+y)/19=y/6 then you get y=6x/13 y'=6x'/13 but remember you are looking for x'+y'

myininaya (myininaya):

does this help?

OpenStudy (anonymous):

yeah...... but what's the final answer.

myininaya (myininaya):

i got 5ft/s+30/13 ft/s

myininaya (myininaya):

[13(5)+30]/13 ft/s=[45+30]/13ft/s=75/13 ft/s